Business Statistics

A sample of 100 mortgages approved during the current year showed that 31 were issued to a single-earner family or individual. The historical average is 25percent. (a) At the .05 level of significance in a right-tailed test, has the percentage of single-earner or individual mortgages risen? Include a sketch of yourdecision rule and show all work. (b) Is this a close decision? (c) State any assumptions that are required.
Answer #1
Given information isA sample of 100(n) mortgages approvedduring the current year showed that 31(X) were issuedtoasingle-earner family or individual. The historical averageis 25(P = 0.25) percent.Sample proportion = p =X/n = 31/100 = 0.31Null Hypothesis: P =0.25Alternative Hypothesis:P>0.25The test statistic isgiven bythe critical value of 'z'at 0.05, the level of significance is 1.645.Since the calculated valueof 'z' is less than the critical vlaue of 'z' so we accept the nullhypothesis and conclude that the percentage of singleearner orindividual mortagages are not risen.b) No it is not a closedecision andc) we assume that thisdata follows normal distribution
answered by: SherylA